The number of hours outside of class that students spend per week on a business statistics course is approximately normally distributed with a mean of 6 hours and a standard deviation 3.2 hours. Suppose there are 300 students registered for this course, approximately how many students work for more than four hours per week?
1
Expert's answer
2020-10-20T18:05:24-0400

"Given \; that, μ=6, σ=3.2, then,\\
a) P(x>4) = P( Z >\frac{4-6}{3.2})\\
=P(Z>-0.63)\\=0.5+ P(0<Z<0.63)=\\
0.5+0.2357=0.7357\\
\text{number of students}=300\times 0.7357\\
=220.71\approx 221\;student\\"